OpenStudy (anonymous):

Find an equation for the line that passes through the points (-2,-4) and (4,-2)

OpenStudy (anonymous):

what are they choices?

OpenStudy (anonymous):

there isnt choices

sam (.sam.):

Use point slope form \(y-y_1=m(x-x_1)\) We need the m, \[m=\frac{-4-(-2)}{-2-4}=\frac{1}{3}\] Choose any of that 2 coordinates, \[y-(-2)=\frac{1}{3}(x-4)\]

OpenStudy (anonymous):

ok, now what?

sam (.sam.):